OpenStudy (anonymous):

Will figure out how to graph on my own but help on what is the Center and what is the radius? x^2+y^2+6x-16y+48=0 would the radius be 0?

OpenStudy (campbell_st):

you need to complete the square in the x terms and y terms so rewriting the equation its \[x^2 + 6x + y^2 - 16y = -48\] so complete the square for x \[x^2 + 6x + ? = (x + 3)^2\] you need a similar process for terms in y... hope this makes sense